# SuperARC: An Agnostic Test for Narrow, General, and Super Intelligence Based On the Principles of Recursive Compression and Algorithmic Probability

### v2

@@ Abstract
We introduce an open-ended test grounded in algorithmic probability that can avoid benchmark contamination in the quantitative evaluation of frontier models in the context of their Artificial General Intelligence (AGI) and Superintelligence (ASI) claims. Unlike other tests, this test does not rely on statistical compression methods (such as GZIP or LZW), which are more closely related to Shannon entropy than to Kolmogorov complexity and are not able to test beyond simple pattern matching. The test challenges aspects of AI, in particular LLMs, related to features of intelligence of fundamental nature such as synthesis and model creation in the context of inverse problems (generating new knowledge from observation). We argue that metrics based on model abstraction and abduction (optimal Bayesian ‘inference’) for predictive ‘planning’ can provide a robust framework for testing intelligence, including natural intelligence (human and animal), narrow AI, AGI, and ASI. We found that LLM model versions tend to be fragile and incremental as a result of memorisation only with progress likely driven by the size of training data. The results were compared with a hybrid neurosymbolic approach that theoretically guarantees universal intelligence based on the principles of algorithmic probability and Kolmogorov complexity. The method outperforms LLMs in a proof-of-concept on short binary sequences. We prove that compression is equivalent and directly proportional to a system’s predictive power and vice versa. That is, if a system can better predict it can better compress, and if it can better compress, then it can better predict. Our findings strengthen the suspicion regarding the fundamental limitations of LLMs, exposing them as systems optimised for the perception of mastery over human language.
Keywords: ARC tests, prediction, compression, program synthesis, inverse problems, symbolic regression, comprehension, Superintelligence, Generative AI, symbolic computation, hybrid computation, Neurosymbolic computation.
@@ 1 Introduction
A common psychological perspective sees intelligence through the lens of IQ tests, particularly the g-factor, a psychometric construct introduced by Spearman that quantifies the positive correlations between cognitive abilities. This framework is consistently linked to a human-centric perspective of what intelligence is and, therefore, biased towards circular reasoning. The concept of intelligence testing has been explored by researchers in different fields, including starting with machine intelligence rather than biological or human intelligence [2, 3, 4, 5, 6].
In Alan Turing’s famous test [7], it was described how an intelligent system could judge the intelligence of another system. This behaviourist approach defines intelligence not by its internal workings but by its external performance in tasks that require reasoning, problem solving, and communication. However, this perspective has been challenged by those who argue that true intelligence requires intrinsic understanding rather than mere imitation of intelligent behaviour. Douglas Hofstadter [8], for example, argued that defining intelligence in some other way than “that which gets the same meaning out of a sequence of symbols as we do” would support the idea of meaning being an inherent property.
Some scholars argue that intelligence can be objectively defined through tests that evaluate specific computational abilities essential to demonstrate intelligent behaviour, rather than trying to define intelligence itself in absolute terms [2, 6, 3, 4]. This perspective shifts the focus from an abstract or philosophical definition to a practical, measurable framework assessing an entity’s capacity for problem-solving, pattern recognition, and adaptive learning within a structured system. In this regard, Gregory Chaitin [9] proposed that formal definitions of intelligence and measures of its various components should be developed in the context of algorithmic complexity in its application to AI. In the same vein, Solomonoff also explicitly proposed assessing AI capabilities using algorithmic probability [10] and is credited with having solved AI with its optimal inference theory [11].
@@
At recent public events, speaking about the foundations of AI and AGI, some leaders in the AI industry have drawn strong parallels between algorithmic complexity, data compression, and AI [16, 17]. Although these terminologies, such as AGI and ASI, are currently loosely defined in the scientific literature, these claims and the current understanding make the connection between LLMs (or any other generative AI), algorithmic complexity, and data compression clearer and more explicit, even calling it fundamental for general and super intelligence, artificial or natural. One idea expressed by Sutskever [16], is that Stochastic Gradient Descent (SGD), a main iterative optimisation algorithm for optimising an objective function used to train models in machine learning (ML) and artificial intelligence (AI), is a practical approximation to finding a computer program that compresses the encoding data in the search space and performs a type of ‘Kolmogorov search’ to find an implicit small computer program embedded in the weights of a ‘soft computer’ or a neural network such as a large Transformer. In a previous work, we successfully explored some of these ideas, proving that we can perform this search on non-differentiable spaces using metrics purely based on algorithmic complexity to search for those programs in model space, making the previously considered fundamental requirement of differentiality redundant [18]. Encoders are effectively (lossy or lossless) compression heuristics and, therefore, deeply connected to algorithmic complexity via compression. Similar ideas are also in evidence in Schmidhuber’s Gödel machines [19] work and Hutter’s AIXI [11] based on Levin’s search [20] and the principles of Algorithmic Probability [21].
@@ 2 Intelligence and Compression
By proposing a formal and more objective definition of intelligence and based on our previous work on computational irreducibility and unpredictability [34], we propose a test for (Super)intelligence based on Algorithmic Information Theory (AIT) [35] specifically testing recently strongly associated features with intelligence in the context of discussions of Artificial General Intelligence (AGI)  [30, 31, 36, 37], such as model abstraction, generalisation, and planning. Here we will argue that all or most of these features are related to just three, therefore, one feature measured by three methods:
@@ 3 Assessing the capabilities of frontier models and LLMs
    LLMs and other AI systems are not self-driven and as such cannot be reasoning agents on their own; they only act upon being triggered and prompted by humans, otherwise they do not posses any internal states (e.g. activity when not prompted).
@@ 4.2.1 Why CTM and BDM as standard for abstraction and planning
A limitation of CTM is that running CTM to approximate model compression and achieve optimal prediction is computationally very expensive. If there were infinite resources, CTM would perform perfect recursive compression and provide the most optimal answer to any computable question given an observation. However, even with access to infinite resources, there are no theoretical or practical guarantees of LLM convergence to any optimal answer. In practice, LLMs are currently more expensive in applications where approaches like CTM could deliver better results (such as for this benchmark, empirically proven to better characterise questions and predict answers encoded in the form of binary sequences) without spending billions of USD in training giant neural systems like LLMs. However, our points is that one does not need to pick one over the other as they can be combined to provide the best approximation to both an optimal but efficient path to an answer under time and resource restrictions. In this regard, CTM/BDM is a resource-bounded approximation to optimal inference that combines pure forms of each side (neuro–based on classical statistics, and symbolic–based on optimal theory). In this sense, the CTM/BDM combo represents the purest form of neurosymbolic computation with no extra steps.
@@ 4.3 Comprehension via Algorithmic Probability
In this paper, the comprehension of LLMs is evaluated using these principles from algorithmic complexity and algorithmic probability. The test is designed to assess the model’s ability to generate code or mathematical models/formulae that compress sequences of increasing complexity. Non-binary sequences are categorized into three levels–Low, Medium, and High complexity–representing datasets that exhibit simple, intricate, and random patterns, respectively. Binary sequences, on the other hand, are classified as either random or ‘climbers’ (as defined in the following section). Thus, a pragmatic compression-as-comprehension test is designed and applied to various LLM models and versions, encompassing test elements of diverse complexity classes which can be understood and compared individually and collectively.
In other words, the SuperARC framework assesses how the LLM model is able to generate an algorithm $\mathcal{A}$ such that, when applied to the input dataset $\tau$ it is able to compress such input by learning its features mechanically and producing a compressed representation $\partial$. Then, by inverting such am algorithm and obtaining the algorithm $\mathcal{A}^{-1}$, the inputs $\tau$ are obtained losslessly with minimal complexity of the combined algorithms according to a complexity metric $\mathcal{M}$. Formally, the LLM is presented with the following task:
@@ 4.4 Design of Experiments
  • Next-digit prediction task with binary and non-binary sequences: We prompted large language models (LLMs) specialising in time series forecasting to predict the digits of non-binary sequences of increasing complexity of two type. The first type are random binary sequences according to increasing CTM/BDM, and the second type are called ‘climbers’. These are strings that when sorted by algorithmic probability in descending order, or algorithmic complexity in ascending order, these binary sequences are longer than strings in their same complexity group (strings with the same or very close complexity values as measured by BDM), hence their complexity is not driven by string length but by their internal structure, aligning with an intuitive understanding of simplicity vs randomness in sequence structure [74]. In other words, these are strings that clearly correspond to lower randomness values because they show lower complexity estimations compared to shorter strings approximated by BDM in the vicinity. For example, the sequence 0101010101… up to certain finite size $n$ is clearly less algorithmic random and therefore more algorithmic probable than any other more random looking string, short or long of the same size $n$ and therefore such a patterned sequence would appear much earlier in a complexity hierarchy and as standing out when sorting strings by both complexity and length. While $01$… is just an example with a clear statistical pattern, CTM is able to detect strings that have non-statistical patterns but are still of very low algorithmic complexity. So, we tested these strings against LLMs capabilities to spot them.
@@ 6 SuperARC-seq
$\delta_{k}=\frac{n_{k}}{\displaystyle\sum_{j=1}^{n_{k}}\delta_{k,j}}\text{ for}R_{k,j}\text{ of type }k,$ (7)
@@ 6.1 Applying SuperARC-seq
It is important to notice that the SuperARC-seq application hereby considered only took into account binary sequences. Whenever integer sequences were considered, a clear biasing of the results was observed as LLMs started to take advantage of their training corpus to actually show memorization rather than abstraction/comprehension. Figures 7 and 8 present the percentages of each output types and test scores when different types os sequences were considered, respectively.
@@
Figure 8: Test scores when different types of sequences are considered. Consistent with the results shown in Figure 7, the inclusion of integer sequences leads to significantly higher test scores for the LLMs. This outcome arises from the models’ ability to exploit their internalized training data by directly associating observed sequences with pre-learnt formulas, suggesting a form of hash-like memorization. These findings highlight the importance of restricting the evaluation to binary sequences in order to obtain an unbiased measure of each model’s true performance, as such sequences are less likely to have been included in the models’ training corpora.
@@ 7 Conclusions
The models’ outputs suggest strength in replication but a lack of adaptive and ‘inventive thinking’. The predominance of trivial or incorrect solutions demonstrates an inability to think ‘outside the box’ (as in if it had not been seen in the training distribution). This suggests that while LLMs can mimic comprehension through retrieval, pattern matching, and Chain-of-Thought techniques, their capabilities remain bounded when tested against algorithmically complex sequences.

### v4

# SuperARC: Can Increasing Complexity Explain Intelligence? A Test for Artificial Super Intelligence Based On the Principles of Causal Recursive Compression and Algorithmic Probability

@@ Abstract
We introduce an open-ended test grounded in Kolmogorov-Chaitin complexity, information theory, and algorithmic probability that can avoid benchmark contamination in the quantitative evaluation of frontier models in the context of their Artificial General Intelligence (AGI) and Superintelligence (ASI) claims. Unlike other tests, this test does not rely on statistical compression methods (such as GZIP or LZW), which are more closely related to Shannon entropy than to Kolmogorov-Chaitin complexity and are not able to test beyond simple pattern matching. The test challenges aspects of AI, in particular LLMs, related to features of intelligence of fundamental nature such as synthesis and model creation in the context of inverse problems (generating new knowledge from observation). We argue that metrics based on model abstraction and abduction (optimal Bayesian ‘inference’) for predictive ‘planning’ can provide a robust framework for testing intelligence, including natural intelligence (human and animal), narrow AI, AGI, and ASI. We found that LLM model versions tend to be fragile and incremental as a result of memorisation only with progress likely driven by the size of training data. The results were compared with a hybrid neurosymbolic approach that theoretically guarantees universal intelligence based on the principles of algorithmic probability and Kolmogorov complexity. The method outperforms LLMs in a proof-of-concept on short binary sequences. We prove that compression is equivalent and directly proportional to a system’s predictive power and vice versa. That is, if a system can better predict it can better compress, and if it can better compress, then it can better predict. Our findings strengthen the suspicion regarding the fundamental limitations of LLMs, exposing them as systems optimised for the perception of mastery over human language.
Keywords: ARC tests, prediction, compression, program synthesis, inverse problems, causal AI, symbolic regression, comprehension, Superintelligence, Generative AI, symbolic computation, hybrid computation, Neurosymbolic computation.
@@ 1 Introduction
A common psychological perspective sees intelligence through the lens of IQ tests, particularly the g-factor, a psychometric construct introduced by Spearman that quantifies the positive correlations between cognitive abilities. This framework is consistently linked to a human-centric perspective of what intelligence is and, therefore, biased towards circular reasoning. The concept of intelligence testing has been explored by researchers in different fields, including starting with machine intelligence rather than biological or human intelligence [2, 3, 4, 5, 6]. Some scholars argue that intelligence can be objectively defined through tests that evaluate specific computational abilities essential to demonstrate intelligent behaviour, rather than trying to define intelligence itself in absolute terms [2, 6, 3, 4]. This perspective shifts the focus from an abstract or philosophical definition to a practical, measurable framework assessing an entity’s capacity for problem-solving, pattern recognition, and adaptive learning within a structured system. It reflects an operational turn in the study of intelligence, emphasizing the design of formal benchmarks and quantifiable metrics. However, this approach is not without its philosophical challenges. By reducing intelligence to observable outputs, it risks overlooking the role of internal representation, consciousness, or semantic understanding–dimensions emphasized in critiques like Searle’s. In response to such concerns, researchers have sought to ground their metrics in more fundamental notions of computation and inference. For example, Gregory Chaitin [7] proposed that formal definitions of intelligence and its components should emerge from the mathematical theory of algorithmic complexity. Similarly, Solomonoff [8] advanced the idea of evaluating intelligence through algorithmic probability, laying the foundation for optimal prediction frameworks. These formal approaches, further developed in universal models like Hutter’s AIXI [9], attempt to reconcile objective evaluation with theoretical generality, but they still provoke debate regarding their ability to fully capture the qualitative essence of intelligence.
@@
At recent public events, speaking about the foundations of AI and AGI, some leaders in the AI industry have drawn strong parallels between algorithmic complexity, data compression, and AI [14, 15]. Although these terminologies, such as AGI and ASI, are currently loosely defined in the scientific literature, these claims and the current understanding make the connection between LLMs (or any other generative AI), algorithmic complexity, and data compression clearer and more explicit, even calling it fundamental for general and super intelligence, artificial or natural. One idea expressed by Sutskever [14], is that Stochastic Gradient Descent (SGD), a main iterative optimisation algorithm for optimising an objective function used to train models in machine learning (ML) and artificial intelligence (AI), is a practical approximation to finding a computer program that compresses the encoding data in the search space and performs a type of ‘Kolmogorov search’ to find an implicit small computer program embedded in the weights of a ‘soft computer’ or a neural network such as a large Transformer. In a previous work, we successfully explored some of these ideas, proving that we can perform this search on non-differentiable spaces using metrics purely based on algorithmic complexity to search for those programs in model space, making the previously considered fundamental requirement of differentiability redundant [16]. Encoders are effectively (lossy or lossless) compression heuristics and, therefore, deeply connected to algorithmic complexity via compression. Similar ideas are also in evidence in Schmidhuber’s Gödel machines [17] work and Hutter’s AIXI [9] based on Levin’s search [18] and the principles of Algorithmic Probability [19].
@@ 2 Intelligence and Compression
By proposing a formal and more objective definition of intelligence and based on our previous work on computational irreducibility and unpredictability [32], we propose a test for (Super)intelligence based on Algorithmic Information Theory (AIT) [33] specifically testing recently strongly associated features with intelligence in the context of discussions of Artificial General Intelligence (AGI) [28, 29, 34, 35, 36, 37, 38, 39, 40]. Here we will argue that all or most of these features are related to just three, therefore, one feature measured by three methods:
@@ 3 Assessing the capabilities of frontier models and LLMs
    LLMs and other AI systems are not self-driven and as such cannot be reasoning agents on their own; they only act upon being triggered and prompted by humans, otherwise they do not possess any internal states (e.g. activity when not prompted).
@@ 4.2.1 Why CTM and BDM as standard for abstraction and planning
A limitation of CTM is that running CTM to approximate model compression and achieve optimal prediction is computationally very expensive. If there were infinite resources, CTM would perform perfect recursive compression and provide the most optimal answer to any computable question given an observation. However, even with access to infinite resources, there are no theoretical or practical guarantees of LLM convergence to any optimal answer. In practice, LLMs are currently more expensive in applications where approaches like CTM could deliver better results (such as for this benchmark, empirically proven to better characterise questions and predict answers encoded in the form of binary sequences) without spending billions of USD in training giant neural systems like LLMs. However, our point is that one does not need to pick one over the other as they can be combined to provide the best approximation to both an optimal but efficient path to an answer under time and resource restrictions. In this regard, CTM/BDM is a resource-bounded approximation to optimal inference that combines pure forms of each side (neuro–based on classical statistics, and symbolic–based on optimal theory). In this sense, the CTM/BDM combo represents the purest form of neurosymbolic computation with no extra steps.
@@ 4.3 Comprehension via Algorithmic Probability
In this paper, the comprehension of LLMs is evaluated using these principles of algorithmic complexity and algorithmic probability. The test is designed to assess the model’s ability to generate code or mathematical models/formulae that compress sequences of increasing complexity. Non-binary sequences are categorised into three levels–Low, Medium, and High Complexity–representing datasets that exhibit simple, intricate, and random patterns, respectively. Binary sequences, on the other hand, are classified as either random or what we call ‘climber’ strings, low complexity strings as defined in the following section. Thus, a pragmatic compression-as-comprehension test is designed and applied to various LLM models and versions, encompassing test elements of diverse complexity classes which can be understood and compared individually and collectively.
In other words, the SuperARC framework assesses how the LLM model is able to generate an algorithm $\mathcal{A}$ such that, when applied to the input data set $\tau$, it is able to compress this input by mechanically learning its features and producing a compressed representation $\partial$. Then, by inverting such am algorithm and obtaining the algorithm $\mathcal{A}^{-1}$, the inputs $\tau$ are obtained losslessly with minimal complexity of the combined algorithms according to a complexity metric $\mathcal{M}$. Formally, the LLM is presented with the following task:
@@ 4.4 Design of Experiments
  • Next-digit prediction task with binary and non-binary sequences: We prompted large language models (LLMs) specialising in time series forecasting to predict the digits of non-binary sequences of increasing complexity of two type. The first type are random binary sequences according to increasing CTM/BDM, and the second type are called ‘climbers’.
    • Climbers are strings that when sorted by algorithmic probability in descending order (highest to lowest probability), or algorithmic complexity in ascending order (lowest to highest randomness), these binary sequences are longer than strings in their same complexity group defined as strings with the same or very close complexity values as measured by BDM but of significantly longer length than them. This means that for these strings, their complexity is definitively not driven by string length only but by (simple) their internal structure, aligning with an intuitive understanding of simplicity vs. randomness in sequence structure [77]. In other words, these are strings that clearly correspond to lower randomness values because they show lower complexity estimations compared to shorter strings in the vicinity. For example, the sequence 0101010101… up to certain finite size $n$ is clearly less algorithmic random and therefore more algorithmic probable than any other more random looking string, short or long of the same size $n$, and therefore such a patterned sequence must appear earlier in a complexity hierarchy if BDM works correctly. So, knowing these are highly structured strings with high algorithmic probability, we tested whether LLMs would identify them by producing short models and better predictions for them compared to others.
@@ 6 SuperARC-seq
$\delta_{k}=\frac{n_{k}}{\displaystyle\sum_{j=1}^{n_{k}}\delta_{k,j}^{-1}}\text{ for }R_{k,j}\text{ of type }k,$ (7)
@@ 6.1 Applying SuperARC-seq
It is important to notice that the SuperARC-seq application hereby considered only took into account binary sequences. Whenever integer sequences were considered, a clear biasing of the results was observed as LLMs started to take advantage of their training corpus to actually show memorisation rather than abstraction/comprehension. Figures 7 and 8 present the percentages of each output types and test scores when different types os sequences were considered, respectively.
@@
Figure 8: Test scores when different types of sequences are considered. Consistent with the results shown in Figure 7, the inclusion of integer sequences leads to significantly higher test scores for the LLMs. This outcome arises from the models’ ability to exploit their internalized training data by directly associating observed sequences with pre-learned formulas, suggesting a form of hash-like memorisation. These findings highlight the importance of restricting the evaluation to binary sequences in order to obtain an unbiased measure of each model’s true performance, as such sequences are less likely to have been included in the models’ training corpora.
@@ 7 Conclusions
The models’ outputs suggest strength in replication but a lack of adaptive and ‘inventive thinking’. The predominance of trivial or incorrect solutions demonstrates an inability to think ‘outside the box’ (as in if it had not been seen in the training distribution). This suggests that while LLMs can mimic comprehension through retrieval, pattern matching, and Chain-of-Thought techniques, their capabilities remain bounded when tested against algorithmically complex sequences. These observations point directly to a key distinction between current systems and Strong AI: the latter would require the ability to autonomously generate new strategies, abstract concepts, and exhibit flexible problem solving beyond training data. In contrast, the limitations seen here highlight how existing LLMs remain confined to narrow intelligence and lack the dynamic reasoning abilities expected of Artificial General Intelligence (AGI).